# latexText page

### v1

@@ 1 Introduction
To perform opponent modeling we assume that we are in the repeated game setting where the game is repeated for some number of iterations. This number may be finite, infinite, or unknown (and potentially infinite). Defining repeated games in the strategic-form setting is relatively straightforward. At each iteration $t$, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the pure strategy selected by the opponent $s^{t}_{-i}$ (though we do not know the opponent’s mixed strategy). We remember all of these strategies for all subsequent iterations. Therefore, at iteration $t$, our strategy can be dependent on $\sigma^{1}_{i},\ldots,\sigma^{t-1}_{i}$ as well as $s^{1}_{-i},\ldots,s^{t-1}_{-i}.$ In perfect-information sequential games, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the pure actions taken by the opponent at nodes along the path of play (we do not know what the opponent would have done at a node that was not reached). Again, at time $t$ our strategy can be dependent on these values for times from 1 to t1.t-1.
@@ 2 Consistent opponent modeling
If we can show that BBR is not consistent in a strategic-form game, this directly implies that it is also not consistent in extensive-form games of perfect and imperfect information. Consider the game of Rock-Paper-Scissors and suppose that $\sigma^{*}_{-i}=(0.8,0.1,0.1)$ for (R,P,S). Suppose that we are using BBR with $k=3$ sampled strategies: (0.5,0.3,0.2), (0.3,0.5,0.2), (0.2,0.3,0.5). BBR choosing the opponent model to play each of the sampled strategies in proportion to their posterior value; therefore, the opponent model is always a convex combination of the sampled strategies. Since $\sigma^{*}_{-i}$ is clearly not close to any convex combination of the sampled strategies, clearly $M_{t}$ cannot approach $\sigma^{*}_{-i}$. ∎
@@ 5 Related research
While we have used a full best response strategy for all our algorithms, note that the meta-algorithm presented as Algorithm 1 allows for arbitrary response function $r_{i}.$ A best response may not be robust if the opponent’s true strategy differs significantly from the opponent model, and other approaches may be preferable. Potentially more robust alternatives include a restricted Nash response [6] and related concept of an $\epsilon$-safe best response [10]. These approaches balance between exploitation and ability, maximizing performance against the opponent’s strategy subject to a constraint on our strategy’s exploitability. Note that these approaches are specific to two-player zero-sum games, while our new algorithm also applies to non-zero-sum games. An alternative approach for achieving robustness may be to apply counterfactual regret minimization (CFR) [12] or one of its sampling variants [9]. CFR maintains counterfactual regret values for each action at each information set, and a popular variant uses regret matching to select the actions in proportion to their positive counterfactual regret at each iteration; this may be more robust than playing a pure best response. Note that CFR is not an alternative to our opponent modeling algorithm; it would just provide an alternative procedure to responding to the opponent model. The application of CFR in real time would require the aid of an opponent modeling algorithm to produce baseline counterfactual values for unobserved paths of play. For example, if we are player 1 in Kuhn poker and bet with a King and player 2 folds, we would not have any information about what our payoff would have been had we chosen to check instead unless we had an opponent model to use to estimate the counterfactual value. All of the alternative response approaches described in this paragraph require an accurate opponent model to be successful.

### v3

# Consistent Opponent Modeling of Static Opponents in Imperfect-Information Games

@@ 1 Introduction
To perform opponent modeling we assume that we are in the repeated game setting where the game is repeated for some number of iterations. This number may be finite, infinite, or unknown (and potentially infinite). Defining repeated games in the strategic-form setting is relatively straightforward. At each iteration $t$, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the vector of pure strategies selected by all players $\mathbf{s^{t}}$ (though of course we do not know opponents’ mixed strategies). We remember all of these strategies for all subsequent iterations. Therefore, at iteration $t$, our strategy can be dependent on $\sigma^{1}_{i},\ldots,\sigma^{t-1}_{i}$ as well as $\mathbf{s^{1}},\ldots,\mathbf{s^{t-1}}$. In perfect-information sequential games, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the actions taken by all players at nodes along the path of play (we do not know what an opposing player would have done at a node that was not reached). Again, at time $t$ our strategy can be dependent on these values for times from 1 to t1.t-1.
@@ 2 Consistent opponent modeling
If we can show that BBR is not consistent in a strategic-form game, this directly implies that it is also not consistent in extensive-form games of perfect and imperfect information. Consider the game of Rock-Paper-Scissors and suppose that $\sigma^{*}_{-i}=(0.8,0.1,0.1)$ for (R,P,S). Suppose that we are using BBR with $k=3$ sampled strategies: (0.5,0.3,0.2), (0.3,0.5,0.2), (0.2,0.3,0.5). BBR chooses the opponent model to play each of the sampled strategies in proportion to their posterior value; therefore, the opponent model is always a convex combination of the sampled strategies. Since $\sigma^{*}_{-i}$ is clearly not close to any convex combination of the sampled strategies, clearly $M_{t}$ cannot approach $\sigma^{*}_{-i}$. ∎
@@ 5 Related research
While we have used a full best response strategy for all our algorithms, note that the meta-algorithm presented as Algorithm 1 allows for arbitrary response function $r_{i}.$ A best response may not be robust if the opponent’s true strategy differs significantly from the opponent model, and other approaches may be preferable. Potentially more robust alternatives include a restricted Nash response [6] and related concept of an $\epsilon$-safe best response [10]. These approaches balance between exploitation and safety, maximizing performance against the opponent’s strategy subject to a constraint on our strategy’s exploitability. Note that these approaches are specific to two-player zero-sum games, while our new algorithm also applies to non-zero-sum games. An alternative approach for achieving robustness may be to apply counterfactual regret minimization (CFR) [12] or one of its sampling variants [9]. CFR maintains counterfactual regret values for each action at each information set, and a popular variant uses regret matching to select the actions in proportion to their positive counterfactual regret at each iteration; this may be more robust than playing a pure best response. Note that CFR is not an alternative to our opponent modeling algorithm; it would just provide an alternative procedure to responding to the opponent model. The application of CFR in real time would require the aid of an opponent modeling algorithm to produce baseline counterfactual values for unobserved paths of play. For example, if we are player 1 in Kuhn poker and bet with a King and player 2 folds, we would not have any information about what our payoff would have been had we chosen to check instead unless we had an opponent model to use to estimate the counterfactual value. All of the alternative response approaches described in this paragraph require an accurate opponent model to be successful.